OpenStudy (anonymous):

Decide whether the equation is a trigonometric identity. Explain your reasoning. cos^2x(1+tan^2x)=1 So what would I do to verify first?

OpenStudy (loser66):

1+ tan^2 = sec^2 cos = 1/sec done

OpenStudy (anonymous):

So what does this mean? Is it an identity? and is that all i have to do?

OpenStudy (loser66):

It means you manipulate the left hand side to get the right hand side.

OpenStudy (anonymous):

is that all?...because i wish i could say that i understand...but I dont...i have no idea how to do precalc!

OpenStudy (loser66):

1+ tan^2 = sec^2 ok?

OpenStudy (anonymous):

okay...ya i get that...so would it turn into secxtanx(1-sin^2x)=sinx ?? beacuse thats what my notes say...?

OpenStudy (loser66):

hence, the left hand side becomes cos^2 (sec^2) , right?

OpenStudy (anonymous):

wouldn't it turn into sin^2x?...oh wait? left hand side? i think?

OpenStudy (loser66):

forget your sin look at my guide, the left hand side becomes cos^2 (sec^2) ok?

OpenStudy (anonymous):

okay...ya..sorry...so it turns into cos^2x(sec^2)...got it

OpenStudy (loser66):

and cos = 1/sec ok?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

that is cos^2 = 1/ sec^2 ok?

OpenStudy (anonymous):

yea i can see how that works...i think i follow

OpenStudy (loser66):